OpenStudy (anonymous):

Solve the polynomial equation x3 + 4x2 - x - 4 = 0 & please show work

OpenStudy (anonymous):

\(x^3+4x^2-x-4=0 \implies x^2(x+4)-(x+4)=0 \implies (x+4)(x^2-1)=0\) \( \implies x=-4 \text{, } x=1 \text{ or }x=-1\).